Suppose we have a group of 50 TV’s and decide to sample 15. What is the probability that the 15 selected contain the best 5 TV’s? 



1
Expert's answer
2021-10-06T16:55:49-0400

It follows that the probability "p" that the 15 selected contain the best 5 TV’s is equal to


"p=\dfrac{{50 \choose 5}}{{50 \choose 15}}=\dfrac{\frac{50!}{5!45!}}{\frac{50!}{15!35!}}=\dfrac{15!35!}{5!45!}
=\dfrac{15\cdot 14\cdot 13\cdot 12\cdot 11\cdot 10\cdot 9\cdot 8\cdot 7\cdot 6}{45\cdot 44\cdot 43\cdot 42\cdot 41\cdot 40\cdot 39\cdot 38\cdot 37\cdot 36}"


"=\dfrac{ 7}{ 43\cdot 41\cdot 38\cdot 37\cdot 3}=\dfrac{7}{7436334}."